OpenStudy (anonymous):

1. a square pyramid with base edge 80 m and slant height 100 m (1 point)44,800 m2 48,000 m2 16,000 m2 22,400 m2 2. a cone with slant height 22 cm and radius 7 cm (1 point)about 484 cm2 about 507 cm2 about 637 cm2 about 1,123 cm2 3. a sphere with radius 12 cm (1 point)about 904 cm2 about 576 cm2 about 452 cm2 about 1,809 cm2

OpenStudy (anonymous):

SA = s^2 + 2sl SA = πr^2 + πrl SA = 4πr^2

OpenStudy (anonymous):

use this formula and you will get the answer XD

OpenStudy (anonymous):

Thanks I understand now, I figured out the answers. :)

OpenStudy (anonymous):

that's good xD